# “Medium” LMs of Code in the Era of LLMs: Lessons From StackOverflow

### v2

@@ 4.3. Task 2: Quality Tag prediction
4.4. Task 3: Closed question prediction
We choose a Kaggle competition dataset for predicting closed questions on SO (Kag, 2013) for this downstream task. The provided dataset contains 140,272 samples. The competition motivates this task by saying that SO is a widely-used platform that provides programmers with high-quality answers to their programming questions on a daily basis. At the time this competition was published, SO received over 6,000 questions every weekday and an automatic content moderation would help maintain post quality. One aspect of moderation involves closing questions that are deemed inappropriate for the platform. At the time this competition was published, around 6% of all new questions ended up being closed for various reasons. The purpose of the competition is to analyze said quality and predict when questions should be closed for effective moderation. They excluded the ‘exact duplicate’ closed reason from the dataset as it depends on previous questions. The competition requires the development of a classifier that can predict whether a question will be closed or not, based on the submitted question and the reason for closure. The categories of classification are ‘open’, ‘not a real question’, ‘off topic’, ‘not constructive’ and ‘too localized’.
@@ 4.4. Task 3: Closed question prediction
4.5. Task 4: Obsolete post detection (New)
Figure 6. Answer 4601538 marked as obsolete on StackOverflow
Given the speed at which technology advances, answers on StackOverflow often become outdated. Outdated answers are any of those that propose solutions that are no longer state-of-the-art. They may offer solutions that are less performant or precise than what is now available, use now-deprecated APIs and libraries (be obsolete), and in some cases propose solutions that create security concerns. For instance, there have been more than 25 versions of TensorFlow (tf, 2021) since its release in 2015, including a major change to its entire compute structure (version 2). This library is also a highly popular subject of questions. Many of those are now years old and involve classes, methods, and parameters that have since been renamed. More generally, over 24 million answers contain code snippets, which software engineers often hope to reuse. When such snippets are obsolete, that reuse may severely reduce the quality of their code.
One might hope that such “outdatedness” is easily discovered by e.g. looking at the date posted, or comments pointing out such issues. However, date is not always a good indicator of obsolence (one type of outdatedness); Zhang et al. (Zhang et al., 2019) found that more than half of obsolete answers were already obsolete at the time of posting. Indeed, ‘year of posting’ only poorly explained outdatedness in our analysis as well. Comments are a useful resource: many posts have comments to indicate obsoletion. However, commenting or editing answers is time-consuming, so many posts with obsolete content go unmarked. Downvoting obsolete answers is ineffective, as posts that were once correct and popular require many downvotes to yield a negative score. SO’s maintainers are aware of this problem and have recently proposed explicitly marking outdated answers as such as shown in Figure 6. However, there are only 9 such answers marked as obsolete by SO at the time this paper was written. We thus propose to fine-tune SOBert to be able to predict these obsolete answers based on a manually curated dataset, which we describe next.
4.5.1. Raw Data Collection
@@
Users often leave comments on a post or answer to indicate obsoletion. We selected answer threads on the basis of the selection criteria from (Zhang et al., 2019). Comments were chosen which had one of the four keywords: deprecated, outdated, obsolete or out of date, while the corresponding question posts did not have these keywords in them. We found 85,586 such samples of comments.
@@
4.5.2. Data Annotation
After applying our heuristic selection criteria described above we obtain some 493K samples from the total database. We manually annotate 1000 random samples out of these, sampling equally from the three categories. The goal of the annotation process is to produce samples consisting solely of the text in an answer or comment on the answer along with a classification of obsolete or otherwise. Answers or comments that either include or refer to obsolete code in any way (e.g., comments pointing out obsolete code in the answer or answers edited to point out they contain outdated solutions) are classified as obsolete, regardless of whether newer techniques or alternative solutions are also provided in the same. This way, models trained on this dataset can be used to highlight posts that should be marked as obsolete using SO’s new system based on existing community input.
The annotation protocol was developed by having two professional programmers annotate a pilot set of answers independently. After completion of the pilot annotation, the differences and difficult-to-annotate cases were discussed and details of such cases were documented. This protocol was improved iteratively by both professionals on other pilot sample sets until consensus was reached about the protocol. The data annotation process relied on the protocol and a set of 10 annotated samples that served as guidelines.
The pilot annotation sets were then discarded and the same samples were re-annotated using the finalized protocol. The inter-annotator agreement of the annotations using the finalized protocol has a Cohen’s kappa of 0.63, which indicates that the agreement level is substantial (Landis and Koch, 1977).
@@ 4.5.2. Data Annotation
Task Model Weighted Accuracy Weighted Precision Recall F1-Score
[Uncaptioned image] NER
BertOverflow
0.69
0.67
0.68
BBUBase
0.80
0.46
0.55
BBULarge
0.56
0.83
0.56
0.63
SOBertBase
0.88 0.93
0.88
0.90
SOBertLarge
0.83
0.86
0.93
0.83
SoftNER
0.78
0.80
0.79
[Uncaptioned image] Quality
BertOverflow
0.72
0.72
0.72
0.72
BBUBase
0.82
0.83
0.82
0.82
BBULarge
0.71
0.72
0.71
0.71
SOBertBase
0.86
0.86
0.86
0.86
SOBertLarge
0.92 0.92 0.92 0.92
[Uncaptioned image] Closed Question
BertOverflow
0.53
0.49
0.53
0.41
BBUBase
0.60
0.56
0.60
0.56
BBULarge
0.53
0.48
0.53
0.43
SOBertBase
0.63 0.62 0.63 0.61
SOBertLarge
0.63
0.61
0.63 0.61
[Uncaptioned image] Obsoletion
BertOverflow
0.17
0.98
0.17
0.26
BBUBase
0.72
0.97
0.72
0.82
BBULarge
0.59
0.97
0.59
0.72
SOBertBase
0.87
0.97
0.87
0.91
SOBertLarge
0.90 0.97 0.90 0.93
Table 1. Performance Metrics for the Models Across the 4 Tasks - Named Entity Recognition (NER), Quality Prediction, Closed Question Prediction, and Obsoletion Detection
5. Results and Discussion
Figure 7. Comparison of F1-score performance of various models on 4 down-stream tasks: (1) NER, (2) quality tag prediction, (3) closed question prediction, and (4) obsolete prediction. The models evaluated include BERT (base and large), SOBert (base and large), BERTOverflow. We include SoftNER for the NER task. We also include ChatGPT@0 and ChatGPT@3, which consist of zero-shot and 3-shot prompting ChatGPT respectively, for closed question prediction, quality tag prediction and obsoletion detection on 250 random samples from the test set. SOBert outperforms the other models on all four tasks.
We fine-tune the SOBert models and the three baseline BERT models on the four downstream tasks described in 4. Figure 7 shows the F1-scores attained by the different models juxtaposed with their respective sizes (note the log-scaling and broken x-axis).
In the NER task, we utilized the test set provided by the authors and observed a significant improvement of +22 points in the F1-score when comparing SOBertBase to BERTOverflow as shown in Table 1. For this task, we did not compute the weighted accuracy of models other than the SOBert models, since the original paper (Tabassum et al., 2020) did not incorporate it in their evaluation metrics. We found that SOBertLarge had a greater improvement of +19 in F1-score over BERTLarge, with SOBertBase coming in close behind with an increase of +18 in F1-score over BERTLarge. Both SOBert models had an increase of about +28 over BERTBase. SOBertLarge also had higher accuracy, precision, and recall values of 0.83, 0.93, and 0.83, respectively, compared to BERTLarge’s scores of 0.56, 0.83, and 0.56. Additionally, when compared to the values reported in the paper for SoftNER, which incorporates a context-independent code token classifier with corpus-level features on top of BERTOverflow to further increase performance, we found that both SOBert models had an improvement of ca. +11 in F1-score over SoftNER.
In the second task quality tag prediction (results in Table 1), we found the SOBertLarge model to be particularly powerful, yielding a +20 point F1-score improvement compared to BERTOverflow. Our smaller, SOBertBase also improved over BERTOverflow’s F1 score by +14 points and by +15 points over BERTLarge.
In the third task (results in Table 1), which involves closed question prediction, we observed an increase of +20 in F1-score between both SOBert models and BERTOverflow. This marks the second task, after NER where our Large model is slightly outperformed by the smaller one – we discuss this phenomenon below. The F1-score for closed question prediction was relatively low for all the models in this task, with the highest F1-score achieved by SOBertBase and SOBertLarge at 0.61. While low, these scores are similar to the results reported in recent studies on the same task (Roy and Singh, 2020). The dataset used in this task contains several features, such as CreationDate, Tags, Reputation, Title, BodyMarkdown, etc. Roy and Singh(Roy and Singh, 2020) have explored the use of feature selection techniques and chose to use BodyMarkdown. We used both the Title and BodyMarkdown.
In the fourth task (results in Table 1) which involves obsoletion detection, we see large gains on the F1-score between SOBert and all the other baselines models. The F1 score of BERTOverflow is significantly lower than that of all other baselines. In 5.2 we delve into potential causes for this behavior and explore possible explanations.
5.1. Relation to BERT
In several experiments, the generic BERT models outperform BERT-Overflow, despite the latter using a very similar architecture and training on StackOverflow data only. In two cases, quality tag and closed question prediction, they even came close to rivaling one of our SOBert models. Their strong performance can likely be attributed by a combination of two causes. Compared to BertOverflow, these BERT models trained on significantly more data (16GB vs. 11GB), which would have helped them learn the patterns in language more effectively. Secondly, the two downstream datasets where they perform best are relatively large, containing 60K and 140K samples respectively. Fine-tuning on a larger dataset allows these models to better overcome the distribution-shift compared to their pre-training setup.
5.2. Relation to BERTOverflow
For the final task of obsoletion detection, we find the F1-score increased by a large margin between both SOBert models and BERTOverflow. Although BERTOverflow converges the fastest during fine-tuning between all 5 models compared for this task, we see that it struggles with this classification task. There could be various factors contributing to unsatisfactory performance, which may include some design decisions taken during pre-training the model such as a small context size and the structure of training data. The SO data BERTOverflow was pre-trained on processed individual sentences in isolation which has lead to a loss of contextual information from the original SO corpus. It is also possible that BERTOverflow was inadequately pre-trained, thus rendering it unable to generalize well to newer data and tasks.
5.3. Relation to model size
In this study, we noticed that both SOBert sizes perform better than both BERT sizes in all of the four tasks we examined. However, we did not find a consistent trend of the larger models (of either kind) outperforming their smaller counterparts. For example, in closed question prediction and obsoletion detection, SOBertBase outperformed SOBertLarge in terms of F1-score. Similarly, BERTLarge did not perform as well as BERTBase in quality tag prediction. Out of the four tasks, the two large models both performed better in exactly half. While this runs counter to the wisdom that larger language models are better (Kaplan et al., 2020), we note that this notion typically applies to models that are not fine-tuned but rather prompted (queried). Fine-tuning adds the noteworthy constraint that larger models represent their inputs in more hidden dimensions, which means that the task-specific prediction layer(s) also require more parameters. For instance, predicting a 1-dimensional label from a 768-dimensional hidden state (for the Base models) requires training half as many new parameters as transforming the 1,536-dimensional states from the Large models. Tuning more parameters also increase the threat of overfitting as models quickly learn to memorize their training inputs on small datasets.
A noteworthy observation from the training process was that, while loss continued to decline (5), downstream task performance saturated around halfway through training, at approximately 50,000 steps, for both SOBertBase and SOBertLarge. While language models generally continue to improve on end tasks as the training loss decreases, this trend does not necessarily hold for models that involve a separate task-specific fine-tuning phase. The additional fine-tuning step appears to enable BERT architectures to reach peak downstream performance well before full convergence of the base pre-trained model.
@@ 5.3. Relation to model size
5.4. Relation to ChatGPT
Finally, we compared the results from SOBertBase and SOBertLarge with ChatGPT on 3 downstream tasks – closed question prediction, question quality prediction and obsoletion. While it would be great to be able to test on more samples, due to limits of testing ChatGPT, we mainly evaluate ChatGPT on a small part of the testing set for each of these tasks. Specifically, we chose 250 random samples from the test data (using test+validation for obsoletion detection, the smallest dataset). We use the GPT3.5 API to get predictions from ChatGPT. We get predictions from both SOBertBase and SOBertLarge on the same 250 samples. We compare this with responses from ChatGPT in 2 settings - with only an instruction and no examples (0-shot prompting) and with three examples drawn from the training data (3-shot prompting). The F1-score results for each of the tasks are shown in Table 2. These results indicate that although ChatGPT can solve many NLP problems quite well, it doesn’t fare too well on tasks in this domain.
A point to consider is that our downstream tasks are essentially classification tasks, and ChatGPT is a decoder-only pre-trained language model as opposed to SOBert which is an encoder-only model. BERT-like models are built upon a bidirectional Masked Language Modeling (MLM) objective that emphasizes the encoding of contextual information. Decoder-only models instead learn to derive representations that are primarily useful for text generation, which prioritizes knowledge from recent tokens. This, combined with its lack of domain-specific fine-tuning (in exchange for much more general capabilities) could explain the lower performance for ChatGPT on these tasks.
Task ChatGPT@0 ChatGPT@3 SOBertBase SOBertLarge
[Uncaptioned image]   Quality
0.39
0.44
0.9
0.89
[Uncaptioned image]   Closed Question
0.16
0.17
0.65
0.64
[Uncaptioned image] Obsoletion
0.56
0.52
0.90
0.88
Table 2. Comparison of task accuracy for SOBertBase, SOBertLarge, ChatGPT@0 and ChatGPT@3.
We started this work by arguing that, while many tasks generally benefit from general purpose LLMs, they do not necessarily guarantee cutting-edge performance for tasks where rich domain-specific data is available. The superior results of even our smallest model, SOBertBase, over both ChatGPT and several general-purpose BERT models on our downstream tasks shows that model size is not the only important factor for performance. Our substantial improvement over BertOverflow further highlights that it is crucial to carefully evaluate the requirements, constraints and data available for a given problem and use an appropriate training design to address the specific challenges at hand.
5.5. Limitations
While our assessment of SOBert includes a comprehensive analysis of four downstream tasks from well-established sources, it is important to recognize that the applicability of our findings to different tasks may differ. We also did not include comparisons with GPT-4 in our study due to its unavailability during our experiments. Additionally, the nature of API access from OpenAI implies that the performance of the language model can fluctuate over time due to model updates without prior notification, potentially affecting its performance. We were also restricted by API limitations for the number of samples we could prompt for and the number of examples provided in the prompt. However, the fact that the LLM is much less accurate in each task implies that it is unlikely to be able to bridge the absence of task specific training with examples alone.
6. Conclusion
We present SOBertBase and SOBertLarge, 125M and 762M parameter models respectively trained exclusively on StackOverflow data. We discuss a number of design decisions inspired by both the training of large language models and the nature of StackOverflow posts that enable us to train our models to a high quality. Among others, we create large samples consisting of entire StackOverflow posts and all their associated comments; we train with large batch sizes using a toolkit that enables efficient parallel training, and train our models extensively. With these configurations, the models we trained could be reproduced with cloud computing resources at a budget of just \$374 and \$1600. We then fine-tuned both our models and an array of BERT baselines on four different downstream tasks: question quality prediction, closed question prediction, named entity recognition and a novel task, obsoletion detection. We compare all our models, as well as ChatGPT (on all but one task) and find that our models consistently achieve top performance in all downstream tasks, with the smaller model regularly performing best. Overall, we underscore the importance of considering the specific needs of a task and the available data when choosing an appropriate model training approach, and provide our guidelines for doing so effectively. Our results highlight that there is ample room for smaller, domain-specific models to outperform general-purpose LLMs at a low budget.
7. Data Availability

### v3

@@ IV-B Task 2: [Uncaptioned image] Quality Tag prediction [19]
IV-C Task 3: [Uncaptioned image]Closed question prediction [20]
We choose a Kaggle competition dataset for predicting closed questions on SO [20] for this downstream task. The provided dataset contains 140,272 samples. The competition motivates this task by saying that SO is a widely-used platform that provides programmers with high-quality answers to their programming questions on a daily basis. At the time this competition was published, SO received over 6,000 questions every weekday and an automatic content moderation would help maintain post quality. One aspect of moderation involves closing questions that are deemed inappropriate for the platform. At the time this competition was published, around 6% of all new questions ended up being closed for various reasons. The purpose of the competition is to analyze said quality and predict when questions should be closed for effective moderation. They excluded the ‘exact duplicate’ closed reason from the dataset as it depends on previous questions. The competition requires the development of a classifier that can predict whether a question will be closed or not, based on the submitted question and the reason for closure. The categories of classification are ‘open’, ‘not a real question’, ‘off topic’, ‘not constructive’ and ‘too localized’.
@@ IV-C Task 3: [Uncaptioned image]Closed question prediction [20]
IV-D Task 4: [Uncaptioned image]Obsolete post detection [Uncaptioned image]
Figure 6: Answer 4601538 marked as obsolete on StackOverflow
Given the speed at which technology advances, answers on StackOverflow often become outdated. Outdated answers are any of those that propose solutions that are no longer state-of-the-art. They may offer solutions that are less performant or precise than what is now available, use now-deprecated APIs and libraries (be obsolete), and in some cases propose solutions that create security concerns. For instance, there have been more than 25 versions of TensorFlow [49] since its release in 2015, including a major change to its entire compute structure (version 2). This library is also a highly popular subject of questions. Many of those are now years old and involve classes, methods, and parameters that have since been renamed. More generally, over 24 million answers contain code snippets, which software engineers often hope to reuse. When such snippets are obsolete, that reuse may severely reduce the quality of their code. One might hope that such “outdatedness” is easily discovered by e.g. looking at the date posted, or comments pointing out such issues. However, date is not always a good indicator of obsolence (one type of outdatedness); Zhang et al. [50] found that more than half of obsolete answers were already obsolete at the time of posting. Indeed, ‘year of posting’ only poorly explained outdatedness in our analysis as well. Comments are a useful resource: many posts have comments to indicate obsoletion. However, commenting or editing answers is time-consuming, so many posts with obsolete content go unmarked. Downvoting obsolete answers is ineffective, as posts that were once correct and popular require many downvotes to yield a negative score. SO’s maintainers are aware of this problem and have recently proposed explicitly marking outdated answers as such as shown in Figure 6. However, there are only 9 such answers marked as obsolete by SO at the time this paper was written. We thus propose to fine-tune SOBert to be able to predict these obsolete answers based on a manually curated dataset, which we describe next.
IV-D1 Raw Data Collection
@@
Users often leave comments on a post or answer to indicate obsoletion. We selected answer threads on the basis of the selection criteria from [50]. Comments were chosen which had one of the four keywords: deprecated, outdated, obsolete or out of date, while the corresponding question posts did not have these keywords in them. We found 85,586 such samples of comments.
@@
IV-D2 Data Annotation
The goal of our annotation process is to produce samples consisting solely of the text in an answer or comment on the answer along with a classification of obsolete or otherwise. We first apply our heuristic selection criteria described above to obtain some 493K samples from the total database. We manually annotate 1000 random samples out of these, sampling equally from the three categories. Answers or comments that either include or refer to obsolete code in any way (e.g., comments pointing out obsolete code in the answer or answers edited to point out they contain outdated solutions) are classified as obsolete, regardless of whether newer techniques or alternative solutions are also provided in the same. This way, models trained on this dataset can be used to highlight posts that should be marked as obsolete using SO’s new system based on existing community input.
The annotation protocol was developed by having two professional programmers annotate a pilot set of answers independently. After completion of the pilot annotation, the differences and difficult-to-annotate cases were discussed and details of such cases were documented. This protocol was improved iteratively by both professionals on other pilot sample sets until consensus was reached about the protocol. The data annotation process relied on the protocol and a set of 10 annotated samples that served as guidelines. The pilot annotation sets were then discarded and the same samples were re-annotated using the finalized protocol. The inter-annotator agreement of the annotations using the finalized protocol has a Cohen’s kappa of 0.63, which indicates that the agreement level is substantial [51].
@@ IV-D2 Data Annotation
V Results
In the NER task, we utilized the test set provided by the authors and observed a significant improvement of +22 points in the F1-score when comparing SOBertBase to BERTOverflow as shown in Table I. We did not compute the balanced accuracy of models other than the SOBert models, since the original paper [15] did not incorporate it in their evaluation metrics. When compared to the values reported in the paper for SoftNER, which incorporates a context-independent code token classifier with corpus-level features on top of BERTOverflow to further increase performance, we found that both SOBert models had an improvement of ca. +11 (SOBertLarge) and +4 (SOBertBase) in F1-score over SoftNER.
In the second task quality tag prediction, we found the SOBertLarge model to be particularly powerful, yielding a +20 point F1-score improvement compared to BERTOverflow. Our smaller, SOBertBase also improved over BERTOverflow’s F1 score by +14 points and by +15 points over BERTLarge.
@@ V Results
Task Model Balanced Accuracy Weighted Precision Recall F1-Score
[Uncaptioned image] NER BertOverflow 0.69 0.67 0.68
BBU 0.80 0.46 0.55
BLU 0.56 0.83 0.56 0.63
SOBertBase 0.88 0.93 0.88 0.90
SOBertLarge 0.83 0.86 0.93 0.83
SoftNER 0.78 0.80 0.79
[Uncaptioned image] Quality BertOverflow 0.72 0.72 0.72 0.72
BBU 0.82 0.83 0.82 0.82
BLU 0.71 0.72 0.71 0.71
GPT3.5@0 0.42 0.41 0.42 0.40
GPT4@0 0.54 0.53 0.54 0.47
GPT3.5@10 0.39 0.39 0.39 0.38
GPT4@10 0.46 0.35 0.34 0.30
SOBertBase 0.86 0.86 0.86 0.86
SOBertLarge 0.92 0.92 0.92 0.92
[Uncaptioned image] Closed Question BertOverflow 0.53 0.49 0.53 0.41
BBU 0.60 0.56 0.60 0.56
BLU 0.53 0.48 0.53 0.43
GPT3.5@0 0.23 0.23 0.23 0.16
GPT4@0 0.35 0.35 0.35 0.29
GPT3.5@10 0.30 0.25 0.30 0.24
GPT4@10 0.41 0.37 0.41 0.37
SOBertBase 0.63 0.62 0.63 0.61
SOBertLarge 0.63 0.61 0.63 0.61
TABLE I: Performance Metrics for the Models Across the 3 Tasks - Named Entity Recognition (NER), Quality Prediction, and Closed Question Prediction
In the third task, which involves closed question prediction, we observed an increase of +20 in F1-score between both SOBert models and BERTOverflow. This marks the second task, after NER where our Large model is slightly outperformed by the smaller one – we discuss this phenomenon below. The F1-score for closed question prediction was relatively low for all the models in this task, with the highest F1-score achieved by SOBertBase and SOBertLarge at 0.61. While low, these scores are similar to the results reported in recent studies on the same task [52]. The dataset used in this task contains several features, such as CreationDate, Tags, Reputation, Title, BodyMarkdown, etc. Roy and Singh[52] have explored the use of feature selection techniques and chose to use BodyMarkdown. We used both the Title and BodyMarkdown.
In the fourth task (results in Table II) which involves obsoletion detection, we see large gains on the F1-score between SOBert and all the other baselines models. While StackLlama’s precision edges out SOBert by a slight margin, it offers substantially lower recall. There is also a substantial difference in their sizes, with StackLlama at 7B parameters being 10 times the size of SOBertLarge. We delve into more details in the following section.
VI Discussion
We started this work by arguing that, while general purpose LLMs perform well on many tasks, they do not guarantee cutting-edge performance for tasks where rich domain-specific data is available, but that this requires careful training of the domain-specific models. The results show that even our smallest model, SOBertBase, concistently outperforms both ChatGPT and several general-purpose BERT models. Our substantial improvement over BertOverflow further highlights that it is crucial to carefully evaluate the requirements, constraints and data available for a given problem when building models in data-constrained domains.
VI-A Relation to BERT
In multiple experiments, the generic BERT models outperform BERT-Overflow despite the latter using a very similar architecture and training on StackOverflow data only. In two cases, quality tag and closed question prediction, they even came close to rivaling one of our SOBert models. Their strong performance can likely be attributed to a combination of two sources. For one, compared to BertOverflow, these BERT models were trained on significantly more data (16GB vs. 11GB), which would have helped them learn the patterns in language more effectively. For another, the two downstream datasets where they perform best are relatively large, containing 60K and 140K samples respectively. Fine-tuning on a larger dataset likely allowed these models to overcome much of the distribution-shift relative to their pre-training setup.
For the final task of obsoletion detection, we find the F1-score increased by a large margin between both SOBert models and BERTOverflow. Although BERTOverflow converges the fastest during fine-tuning between all BERT models we compared for this task, it struggles with this classification task. The structure of its training data is likely a core factor in its performance. During pretraining, it processed SO data as individual sentences in isolation. That removes much of the valuable context from their original SO post. Evidently, the representations learned this way depress performance even after fine-tuning with complete posts.
VI-B Relation to model size
In this study, we noticed that both SOBert sizes perform better than both BERT sizes in all of the four tasks we examined. However, we did not find a consistent trend of the larger models (of either kind) outperforming their smaller counterparts. For example, in closed question prediction and obsoletion detection, SOBertBase outperformed SOBertLarge in terms of F1-score. Similarly, BLU did not perform as well as BBU in quality tag prediction. Out of the four tasks, the two large models both performed better in exactly half.
Method Balanced Accuracy Precision Recall F1
Zero Shot Prompting
ChatGPT3.5 0.72 0.53 0.72 0.49
ChatGPT4 0.59 0.52 0.59 0.50
CodeLlama-7b-Instruct-hf 0.42 0.33 0.28 0.13
StackLlama 0.40 0.36 0.26 0.06
Prompt Engineering
ChatGPT3.5@10+CoT 0.78 0.55 0.78 0.57
ChatGPT4@10+CoT 0.59 0.51 0.59 0.50
Fine-tuning
StackLlama 0.72 0.85 0.72 0.77
CodeLlama-7b-Instruct-hf 0.76 0.83 0.76 0.79
Pre-trained BERT
BertOverflow 0.49 0.47 0.49 0.48
BBU 0.70 0.70 0.70 0.70
BLU 0.48 0.47 0.48 0.47
SOBertBase 0.92 0.76 0.92 0.82
SOBertLarge 0.71 0.81 0.71 0.75
TABLE II: Performance Metrics for the Models for Obsoletion Detection [Uncaptioned image]
While this runs counter to the wisdom that larger language models are better [16], we note that this notion typically applies to models that are not fine-tuned but rather prompted (queried). Fine-tuning adds the noteworthy constraint that larger models represent their inputs in more hidden dimensions, which means that the task-specific prediction layer(s) also require more parameters. For instance, predicting a 1-dimensional label from a 768-dimensional hidden state (for the Base models) requires training half as many new parameters as transforming the 1,536-dimensional states from the Large models. Tuning more parameters also increases the risk of overfitting as models can quickly memorize small datasets.
A noteworthy observation from the training process was that, while loss continued to decline (Figure 5), downstream task performance saturated around halfway through training, at approximately 50,000 steps, for both SOBertBase and SOBertLarge. While language models generally continue to improve on end tasks as the training loss decreases, this trend does not necessarily hold for models that involve a separate task-specific fine-tuning phase. The additional fine-tuning step appears to enable BERT architectures to reach peak downstream performance well before full convergence of the base pre-trained model.
@@ VI-B Relation to model size
VI-C Relation to LLMs
The results indicate that, while LLMs can solve many NLP problems quite well, they do not fare well on the tasks studied in this work. A potential contributing factor is that our downstream tasks are all classification tasks, whereas nearly all LLMs are decoder-only, causal language models, which lend themselves especially to generative tasks. BERT-like models, on the other hand, are built upon a bidirectional Masked Language Modeling (MLM) objective that emphasizes the encoding of contextual information. Another, perhaps more significant factor is the domain-specific and challenging nature of the tasks studied here. Predicting question quality, for instance, is quite challenging even for a human reader (for some examples, see the few-shot prompt in the appendix). Mastering this task evidently requires exposure to a large number of labeled examples. Even 10-shot prompting GPT-4 does not meaningfully shrink the performance gap with SOBert, which sees 60K training examples in each epoch. While LLMs can be fine-tuned, it is impractical to do so for every downstream task. As the results with Llama models in Table II show, this also does not necessarily bridge the gap with purpose-tuned bidirectional models.
One might think that waiting for a newer and bigger LLM to be released could improve performance. However, we see that GPT-4 shows lower performance compared to GPT-3.5 in our experiments, suggesting that simply relying on the next generation of LLMs may not be the solution for such domain-specific tasks. These findings underscore the importance of carefully considering the task at hand and the available data before choosing the appropriate model, rather than defaulting to the latest LLM.
VII Conclusion
In this study, we show that small, domain-specific language models built by combining SE insights with ML best-practices can yield superior results to generalist LLMs and BERT models. We trained SOBertBase and SOBertLarge on SO data with hyper-parameter configurations tailored towards the data characteristics of SO posts. With 125M and 762M parameters respectively, these models were trained with a modest budget of \$374 and \$1600. Our models consistently outperformed baseline models across four software engineering-related tasks: question quality prediction, closed question prediction, named entity recognition, and obsoletion detection. We introduce obsoletion detection as a new benchmark task. We compare our models with a range of other BERT models and with LLMs such as GPT-3.5, GPT-4, CodeLlama, and StackLlama. Despite their comparatively smaller size, the SOBert models consistently outperformed all other models, emphasizing the importance of task-specific model training. Our study demonstrates that for certain tasks, developing tailored, domain-specific models can be a powerful and cost-effective alternative to relying solely on large, general-purpose language models.
VIII DATA AVAILABILITY